If yı(t) and y2(t) are two solutions of the differential equation y2 – y +y = 0, then for any constants cı and c2, cıyı(t) +true or false
Answer #1

false 19 Let y, and I are two solution of y2 + y - ya y² + y - y = out - Yo = 0 CY, + Coq,, C1, C, ER. 22 7W- W= (6,8, + G